# Detection of ChatGPT Fake Science with the xFakeSci Learning Algorithm

### v3

@@ Abstract
ChatGPT and generative AI tools are becoming the new reality. This work is motivated by the premise that “ChatGPT content may exhibit a distinctive behavior that can be separated from scientific articles”. In this study, we demonstrate how we tested this premise in two phases and prove its validity. Subsequently, we introduce xFakeSci, a novel learning algorithm, that is capable of distinguishing ChatGPT-generated articles from publications produced by scientists. The algorithm is trained using network models driven from multiple types of data sources, such as ChatGPT-generated documents achieved by means of prompt-engineering, and PubMed articles. To mitigate over-fitting issues, we incorporate a calibration step that is built upon data-driven heuristics, including ratios. We evaluate the algorithm across multiple datasets covering publication periods and diseases (cancer, depression, and Alzheimer’s). Further, we show how the algorithm is benchmarked against the state-of-the-art (SOTA) algorithms. While the xFakeSci algorithm achieve F1 score ranging from 80% - 94%, SOTA algorithms score F1 values between 38% - 52%. We attribute the noticeable difference to the introduction of calibration and a proximity distance heuristic, which we underscore this promising performance. Indeed, the prediction of fake science generated by ChatGPT presents a considerable challenge. Nonetheless, the introduction of xFakeSci algorithm is a significant step on the way to combating fake science.
@@ 1 Highlights
  • Presenting xFakeSci, a network-driven learning algorithm designed tailored to discern the authenticity of biomedical articles. The algorithm is trained using textual data, and calibrated using data-driven heuristics encompassing factors such as ratios and proximity metrics.
  • Evaluating and benchmarking the algorithm using the state-of-the-art (SOTA) algorithms (e.g., NaiveBayes, Support Vector Machines). Strikingly, the xFakeSci algorithm demonstrates a notable disparity in performance compared to existing SOTA approaches.
@@ 2 Introduction
Here, we briefly demonstrate how the emergence of ChatGPT (alongside many other generative AI tools) has impacted our society today: (1) the launch of many special issues and themes to study, assess, analyze, and test the impact and potential of ChatGPT [17, 18, 19, 20, 21], (2) the adoption of new policies by journals regarding ChatGPT authorship [22, 23, 24, 25, 26, 16], (3) the development of ChatGPT plugins and inclusion in professional services such as Expedia and Slack [27], (4) the creation of educational tools (e.g., Wolfram) and the potential development of learning and educational support tools (e.g., Medical Licensing Examination [28]).
@@ 2.1 Literature Background
Due to the significance of fake science and the imperative for plagiarism detection mechanisms, numerous researchers have investigated this research and employed diverse methodologies. Here, we list some of the most relevant works.
Chaka addressed the detection of content generated by tools such as ChatGPT and assessed it in five different languages. The work used a prompt-engineering approach to generate content in languages including Spanish, French, and German. The research concluded that detecting machine-generated content poses a significant challenge, and further investigations are much-needed to protect against plagiarism [29].
@@
Elkhatat et al. evaluated 15 paragraphs generated by ChatGPT (versions 3.5 and 4.0) using means of other automatic detection tools such as Copyleaks [32] and CrossPlag [33]. The authors concluded that AI detection tools were able to predict ChatGPT-generated content in isolation without mixing with other content. However, when the generated content was perturbed using human-written responses, all detection tools failed to produce accurate or consistent results [34]. Clearly, the xFakeSci algorithm does not suffer from such a deficiency in identifying real articles when mixed with ChatGPT-generated content.
@@
With this introduction, we respond to the urgent call to combat fake science by presenting the xFakeSci algorithm. The algorithm is primarily designed to detect ChatGPT-generated content and distinguish it from real PubMed abstracts of published articles. The algorithm operates in two modes: (a) a Single mode, where it is trained from a single source to predict one class, and (b) Multi-mode where the algorithm processes various types of resources to predict the correct label for each class. As a network-driven algorithm, it is trained by a model that utilizes the largest connected components (LCC) as an admissible heuristic. To mitigate overfitting issues, we implemented a data-driven calibration step that ensures the accurate prediction of documents, utilizing a small sample of training articles (100 articles from each source). In the Methods section, we provide detailed explanations of the following components: (1) the prompt-engineering process of how fake documents were generated in three different diseases using ChatGPT, (2) the phases of testing the premise that content generated from ChatGPT may exhibit certain characteristics that reveal its identity and make it distinguishable from real science, (3) the xFakeSci algorithm, which we designed to predict the class of a given document and determine whether it is real or fake. We outline the computational steps, including constructing the network training models, calibrating using data-driven heuristics, testing the algorithm in multiple contexts (diseases) and data from various publication periods, and finally, (4) benchmarking it against state-of-the-art (SOTA) algorithms as a final step.
@@ 3.1 Outcome of Evaluating the Premise
As mentioned earlier, we investigated the premise that of whether ChatGPT-generated content may exhibit unique characteristics that differ from those observed in scientific articles. We tested this intuition two stages:
@@ 3.1.2 Phase II: Further Testing the Distinctive Behavior in ChatGPT-Generated Documents
To further investigate the premise, we conducted a test to analyze the mean ratios of contributing bigrams extracted from k-Folds against the document word count. This analysis aimed to establish a baseline for assessing the contribution of bigrams to the overall content structure. The results revealed a consistent pattern across all three disease datasets. Specifically, ChatGPT-generated datasets exhibited significantly higher ratios than their scientific publication counterparts in each of the k-Folds used. For instance, in the Alzheimer’s disease dataset, ChatGPT scores were (0.27, 0.30, 0.30, 0.28, 0.28, 0.29), while scientific publications from 2010-2014 scored (0.16, 0.17, 0.16, 0.16, 0.17, 0.16), for 2015-2019 (0.15, 0.16, 0.15, 0.16, 0.14, 0.15), and for 2020-2024 (0.15, 0.15, 0.14, 0.15, 0.14, 0.14). These findings are consistent across the other two diseases, as evident in Table 3. Figures 3 and 3 clearly demonstrate that the k-Folds ratios calculated from ChatGPT-generated data are significantly higher than those derived from scientific publications across different years and scopes. They further illustrate a similar pattern for cancer and depression datasets. This evidence reinforces the notion that ChatGPT-generated content may exhibit distinct characteristics compared to scientific articles.
@@ 3.2 Outcome of Label Prediction of Multi-Mode Classification Experiments
To establish confidence in our methods and ensure consistent performance of the xFakeSci algorithm, we conducted two types of experiments in the subject area of three different diseases. Additionally, we performed experiments to evaluate whether the year of publication plays a factor in class prediction. This section presents the outcomes of experiments utilizing ChatGPT-generated text obtained algorithmically using ChatGPT prompt-engineering, as outlined in Algorithm 1, and scientific publications retrieved from the PubMed web portal [37] related to Alzheimer’s, cancer, and depression diseases.
@@
The table demonstrates that xFakeSci detected all 50 PubMed publications for each disease (TP=50). Additionally, we observed that the algorithm identified the ChatGPT-generated documents to varying extents (TN=25, 41, 38) for depression, cancer, and Alzheimer’s, respectively. It remains concerning that ChatGPT is classified as PubMed with (FP=25), indicating that 50% of the test documents are misclassified as real publications. Further research is needed to investigate and improve performance.
@@ 3.3 Outcome of Publication Period as a Factor for the Multi-Mode Classification Experiments
The purpose of this experiment is to test whether the publication period is a factor in making predictions and assigning labels to a dataset of documents with mixed classes (ChatGPT vs PubMed articles). Here, we use the F1 metric as a measure to present our results and explain the associated performance of our algorithm (captured in Table 6). For each disease dataset extracted from three periods (2020-2024; 2015-2019; 2010-2014), we computed the F1 score.
For the cancer disease dataset, the F1 scores recorded were 91%, 92%, and 94% for the three various periods. These scores show a consistent improvement in predicting the labels with older publication datasets, from 2010 to the present.
@@
For the Alzheimer’s disease dataset, the scores showed a slight improvement from 89% (in 2020-2024) to 90% in (2015-2019), but it dropped back to 89% for the (2010-2014) disease dataset. While the pattern of prediction improvement did not hold as we analyzed older publications, the score did not degrade below the F1 score of 89% for (2020-2024).
3.4 Outcome of xFakeSci Performance Analysis against the SOTA Algorithms
@@
Each algorithm was used as a blackbox and received input from two sources (training and test), and in turn, produced a detailed analysis in the form of (TP, TN, FP, FN). Using these metrics, the F1 score was computed accordingly. Figure 6 visually depicts the F1 scores observed over time for the 5 algorithms (including xFakeSci) in three different diseases, presented by sub-figures, one for each disease. Each sub-figure shows stacked bars for the three periods of publications, and each bar represents the F1 score resulting from a given algorithm. In Table 6, we present the performance metrics, specifically the F1 scores, across different publication periods, starting from the most recent. In the period spanning 2020 to 2024, the F1 scores achieved by xFakeSci ranged from 80% to 91%, while those of the state-of-the-art (SOTA) algorithms fluctuated between 43% and 52%. Moving to the period of 2014 to 2019, xFakeSci demonstrated F1 scores ranging from 80% to 92%, compared to the SOTA algorithms which exhibited F1 scores in the range of 43% to 51%. Lastly, during the period of 2010 to 2014, the F1 scores achieved by xFakeSci fluctuated between 80% and 94%, whereas the F1 scores of the SOTA algorithms were recorded between 38% and 52%. Figure 5 shows a screenshot providing evidence of achieving 91% accuracy measured by the F1 metric, while the F1 scores of the SOTA algorithms fluctuated from 43% to 51%. All three sub-figures show a consistent pattern where xFakeSci clearly outperforms all the other four algorithms. The F1 score is calculated using the following equation 1, and the analysis was done using the scikit-learn library [39].
@@ 4.1 Data Collection
We compiled two distinct types of datasets for the study: (1) Literature dataset: To establish a baseline for comparison and train the xFakeSci algorithm, we utilized the PubMed archive to retrieve scientific articles. We employed three search queries: (a) “Alzheimer’s disease and co-morbidities,” (b) “cancer and co-morbidities,” and (c) “depression.” To assess the influence of publication year on fake science detection, we conducted these searches at five-year intervals, resulting in three distinct datasets for each disease (2010-2024). (2) ChatGPT-generated dataset: We obtained this dataset by programmatically prompting the ChatGPT API (version: 3.5, model: “gpt-3.5-turbo-16k”) to generate simulated articles. The prompt-engineering process is described in detail in the “Methods” section.
The Methods section of this paper comprises two primary components:
  • Prompt-Engineering ChatGPT for Simulated Article Generation: We employed ChatGPT, a generative AI tool, for generating simulated articles. We implemented a prompt-engineering technique to guide the process of text generation in three distinct diseases: Alzheimer’s, cancer, and depression.
@@
4.2 ChatGPT Prompt-Engineering for Article Generation
We utilized the ChatGPT API (version: 3.5, model: “gpt-3.5-turbo-16k”) to engineer prompts for generating simulated articles in the subject areas of real published articles (Alzheimer’s, cancer, and depression). These prompts were parameterized using information from the real articles (search keywords used for article retrieval and the average number of words) to make them comparable to real abstracts. They included three key elements: (a) the role of the prompt: “a biomedical researcher,” (b) the request example: “Generate a list of 20 simulated PubMed-style abstracts,” (c) topic example: “the Alzheimer’s disease,” and (d) specifications: each article must contain ID, Title, and Abstract fields. We also instructed the prompt to generate a valid JSON response with these specifications. The number of words helped offset any bias and made the fake articles comparable to the precise level of detail required (the 200-250 words range is a common requirement by many prominent biomedical informatics journals).
@@ 4.2 ChatGPT Prompt-Engineering for Article Generation
Algorithm 1 The Computational Process of ChatGPT Prompt-Engineering for Article Generation
@@ 4.3 Prediction of Fake Articles using xFakeSci
The xFakeSci algorithm is a network-driven label prediction algorithm designed to distinguish between real scientific articles and machine-generated ones. This entails that the algorithm has two main tasks: training the model and testing it to detect the label of entirely new documents that have never been seen before. In this section, we introduce the computational steps that describe the prediction process, starting with: (1) the construction of network training models, (2) the calibration of the algorithm, and (3) the label prediction for each of the ChatGPT-generated articles.

### v4

@@ Abstract
Generative AI tools exemplified by ChatGPT are becoming a new reality. This study is motivated by the premise that “AI generated content may exhibit a distinctive behavior that can be separated from scientific articles”. In this study, we show how articles can be generated using means of prompt engineering for various diseases and conditions. We then show how we tested this premise in two phases and prove its validity. Subsequently, we introduce xFakeSci, a novel learning algorithm, that is capable of distinguishing ChatGPT-generated articles from publications produced by scientists. The algorithm is trained using network models driven from both sources. To mitigate overfitting issues, we incorporated a calibration step that is built upon data-driven heuristics, including proximity and ratios. Specifically, from a total of a 3952 fake articles for three different medical conditions, the algorithm was trained using only 100 articles, but calibrated using folds of 100 articles. As for the classification step, it was performed using 300 articles per condition. The actual label steps took place against an equal mix of 50 generated articles and 50 authentic PubMed abstracts. The testing also spanned publication periods from 2010 to 2024 and encompassed research on three distinct diseases: cancer, depression, and Alzheimer’s. Further, we evaluated the accuracy of the xFakeSci algorithm against some of the classical data mining algorithms (e.g., Support Vector Machines, Regression, and Naive Bayes). The xFakeSci algorithm achieved F1 scores ranging from 80% to 94%, outperforming common data mining algorithms, which scored F1 values between 38% and 52%. We attribute the noticeable difference to the introduction of calibration and a proximity distance heuristic, which underscores this promising performance. Indeed, the prediction of fake science generated by ChatGPT presents a considerable challenge. Nonetheless, the introduction of the xFakeSci algorithm is a significant step on the way to combating fake science.
@@ 1 Highlights
  • Presenting xFakeSci, a network-driven learning algorithm tailored to discern the authenticity of biomedical articles. The algorithm is trained using textual data, and calibrated using data-driven heuristics encompassing factors such as ratios and proximity metrics.
  • Evaluating and benchmarking the algorithm using some of the most common classical algorithms (e.g., NaiveBayes and Support Vector Machines) as a baseline. Strikingly, the xFakeSci algorithm demonstrates a notable disparity in performance compared to existing classical approaches.
@@ 2 Introduction
In this article, we demonstrate how the emergence of ChatGPT (alongside many other generative AI tools) has impacted our society today: (1) the launch of many special issues and themes to study, assess, analyze, and test the impact and potential of ChatGPT [17, 18, 19, 20, 21], (2) the adoption of new policies by journals regarding ChatGPT authorship [22, 23, 24, 25, 26, 16], (3) the development of ChatGPT plugins and inclusion in professional services such as Expedia and Slack [27], and (4) the creation of educational tools (e.g., Wolfram) and the potential development of learning and educational support tools (e.g., Medical Licensing Examination [28]).
@@ 2.1 Literature Background
Due to the significance of fake science and the imperative for plagiarism detection mechanisms, numerous researchers have investigated fake science detection and employed diverse methodologies. Here, we list some of the most relevant works.
Chaka addressed the detection of content generated by tools such as ChatGPT and assessed it in five different languages. The work used a prompt-engineering approach to generate content in languages including Spanish, French, and German. The research concluded that detecting machine-generated content poses a significant challenge, and further investigations are much needed to protect against plagiarism [29].
@@
Elkhatat et al. evaluated 15 paragraphs generated by ChatGPT (versions 3.5 and 4.0) using means of other automatic detection tools such as Copyleaks [32] and CrossPlag [33]. The authors concluded that AI detection tools were able to predict ChatGPT-generated content in isolation without mixing with other content. However, when the generated content was perturbed using human-written responses, all detection tools failed to produce accurate or consistent results [34].
@@
With this study, we respond to the urgent call to combat fake science by presenting the xFakeSci algorithm. The algorithm is primarily designed to detect ChatGPT-generated content and distinguish it from real PubMed abstracts of published articles. The algorithm operates in two modes: (a) a Single mode, where it is trained from a single source to predict one class, and (b) a Multi mode where the algorithm processes various types of resources to predict the correct label for each class. As a network-driven algorithm, it is trained by a model that utilizes the largest connected components (LCC) as an admissible heuristic. To mitigate overfitting issues, we implemented a data-driven calibration step that ensures the accurate prediction of documents, utilizing a small sample of training articles (100 articles from each source). In the Methods section, we will provide detailed explanations of the following components: (1) the prompt-engineering process of how fake documents were generated in three different diseases using ChatGPT, (2) the phases of testing the premise that content generated from ChatGPT may exhibit certain characteristics that reveal its identity and make it distinguishable from real science, (3) the xFakeSci algorithm, which we designed to predict the class of a given document and determine whether it is real or fake. We outline the computational steps, including constructing the network training models, calibrating using data-driven heuristics, testing the algorithm in multiple contexts (diseases) and data from various publication periods, and finally, (4) benchmarking it against the most common classical data mining algorithms as a verification step.
@@ 3.1 Outcome of Evaluating the Premise
As mentioned earlier, we investigated the premise that of whether AI-generated content may exhibit unique characteristics that differ from those observed in scientific articles. We tested this intuition two stages:
@@ 3.1.2 Phase II: Further Testing the Distinctive Behavior in ChatGPT-Generated Documents
To further investigate the premise, we conducted a test to analyze the mean ratios of contributing bigrams extracted from k-Folds against the document word count. This analysis aimed to establish a baseline for assessing the contribution of bigrams to the overall content structure. The results revealed a consistent pattern across all three disease datasets. Specifically, ChatGPT-generated datasets exhibited significantly higher ratios than their scientific publication counterparts in each of the k-Folds used. For instance, in the Alzheimer’s disease dataset, ChatGPT scores were (0.27, 0.30, 0.30, 0.28, 0.28, 0.29), while scientific publications from 2010-2014 scored (0.16, 0.17, 0.16, 0.16, 0.17, 0.16), for 2015-2019 (0.15, 0.16, 0.15, 0.16, 0.14, 0.15), and for 2020-2024 (0.15, 0.15, 0.14, 0.15, 0.14, 0.14). These findings are consistent across the other two diseases, as evident in Table 4. Figures 3 and 3 clearly demonstrate that the k-Folds ratios calculated from ChatGPT-generated data are significantly higher than those derived from scientific publications across different years and scopes. They further illustrate a similar pattern for the cancer and depression datasets. This evidence reinforces the notion that ChatGPT-generated content may exhibit distinct characteristics compared to scientific articles.
@@ 3.2 Outcome of Label Prediction of Multi-Mode Classification Experiments
To establish confidence in our method and ensure consistent performance of the xFakeSci algorithm, we conducted two types of experiments in the subject area of three different diseases. Additionally, we performed experiments to evaluate whether the year of publication plays a role in class prediction. This section presents the outcomes of experiments utilizing ChatGPT-generated text obtained algorithmically using ChatGPT prompt-engineering, as outlined in Algorithm 1, and scientific publications retrieved from the PubMed web portal [37] related to the Alzheimer’s, cancer, and depression diseases.
@@
Table 6 demonstrates that xFakeSci detected all 50 PubMed publications for each disease (TP=50). Additionally, we observed that the algorithm identified the ChatGPT-generated documents to varying extents (TN=25, 41, 38) for depression, cancer, and Alzheimer’s, respectively. It remains concerning that ChatGPT is classified as PubMed with (FP=25), indicating that 50% of the test documents are misclassified as real publications. Further research is needed to investigate and improve performance.
@@ 3.3 Outcome of Publication Period as a Factor for the Multi-Mode Classification Experiments
The purpose of this section is to test whether the publication period is a factor in making predictions and assigning labels to a dataset of documents with mixed classes (ChatGPT vs PubMed articles). Here, we use the F1 metric as a measure to present our results and explain the associated performance of our algorithm (captured in Table 8). For each disease dataset extracted from three periods (2020-2024; 2015-2019; 2010-2014), we computed the F1 score.
For the cancer disease dataset, the F1 scores recorded were 91%, 92%, and 94% for the three different periods. These scores show a consistent improvement in predicting the labels with older publication datasets, from 2010 to the present.
@@
For the Alzheimer’s disease dataset, the scores showed a slight improvement from 89% (in 2020-2024) to 90% in (2015-2019), but it dropped back to 89% for the (2010-2014) disease dataset. While the pattern of prediction improvements did not hold as we analyzed older publications, the score did not degrade below the F1 score of 89% for (2020-2024).
3.4 Outcome of xFakeSci Performance Analysis against other Data Mining Algorithms
@@
Each algorithm was used as a blackbox and received input from two sources (training and test), and in turn, produced a detailed analysis in the form of (TP, TN, FP, FN). Using these metrics, the F1 score was computed accordingly. Figure 6 visually depicts the F1 scores observed over time for the 5 algorithms (including xFakeSci) in three different diseases, presented by sub-figures, one for each disease. Each sub-figure shows stacked bars for the three periods of publications, and each bar represents the F1 score resulting from a given algorithm. Table 7 captures the performance analysis of the xFakeSci algorithm against other classical data mining algorithms (Naive Bayes, Linear SVM, Classical SVM, and Logistic Regression). The performance is presented using the F1 metric for publications spanning between 2020 and 2024. The table shows that xFakeSci scores ranged from 80% to 91%, while those of the data mining algorithms fluctuated between 43% and 52%.
Moving to the period of 2014 to 2019, xFakeSci demonstrated F1 scores ranging from 80% to 92%, compared to other data mining algorithms which exhibited F1 scores in the range of 43% to 51%. Lastly, during the period of 2010 to 2014, the F1 scores achieved by xFakeSci fluctuated between 80% and 94%, whereas the F1 scores of the other data mining algorithms were recorded between 38% and 52% as shown in Table 8. Figure 5 shows a screenshot providing evidence of achieving 91% accuracy measured by the F1 metric, while the F1 scores of the other data mining algorithms fluctuated from 43% to 51%. All three sub-figures show a consistent pattern where xFakeSci clearly outperforms all the other four algorithms. The F1 score is calculated using Equation 1, and the analysis was done using the scikit-learn library [39].
@@ 4.1 Data Collection
We compiled two distinct types of datasets for the study: (1) Literature dataset: To establish a baseline for comparison and train the xFakeSci algorithm, we utilized the PubMed archive to retrieve scientific articles. We employed three search queries: (a) “Alzheimer’s disease and co-morbidities,” which we generated 1196 JSON records, (b) “cancer and co-morbidities”, generated 1243 JSON records, and (c) “depression” which generated 1513 JSON records. To assess the influence of publication year on fake science detection, we conducted these searches at five-year intervals, resulting in three distinct datasets for each disease (2010-2024). (2) ChatGPT-generated dataset: We obtained this dataset by programmatically prompting the ChatGPT API (version: 3.5, model: “gpt-3.5-turbo-16k”) to generate simulated articles. The prompt-engineering process comprises two primary components:
  • Prompt Engineering ChatGPT for Simulated Article Generation: We employed ChatGPT, a generative AI tool, for generating simulated articles. We implemented a prompt-engineering technique to guide the process of text generation in three distinct diseases: Alzheimer’s, cancer, and depression.
@@
4.2 ChatGPT Prompt Engineering for Article Generation
We utilized the ChatGPT API (version: 3.5, model: “gpt-3.5-turbo-16k”) to engineer prompts for generating simulated articles in the subject areas of real published articles (Alzheimer’s, cancer, and depression). These prompts were parameterized using information from the real articles (search keywords used for article retrieval and the average number of words) to make them comparable to real abstracts. They included three key elements: (a) the role of the prompt: “a biomedical researcher,” (b) the request example: “Generate a list of 20 simulated PubMed-style abstracts,” (c) topic example: “the Alzheimer’s disease,” and (d) specifications: each article must contain ID, Title, and Abstract fields. We also instructed prompts to generate a valid JSON response with these specifications. The number of words helped offset any bias and made the fake articles comparable to the precise level of detail required (the 200-250 words range is a common requirement by many prominent biomedical informatics journals). Table 1 captures the search queries and the number of fake articles generated in the JSON format.
Table 1: Result Counts of Prompt Engineering
Search Query Count (JSON Records)
Alzheimer’s disease and co-morbidities 1196
Cancer and co-morbidities 1243
Depression 1513
Total number of generated articles 3952
@@ 4.2 ChatGPT Prompt Engineering for Article Generation
Algorithm 1 The Computational Process of ChatGPT Prompt Engineering for Article Generation
@@ 4.3 Prediction of Fake Articles using xFakeSci
The xFakeSci algorithm is a network-driven label prediction algorithm designed to distinguish between real scientific articles and machine-generated ones. This entails that the algorithm has two main tasks: training the model and testing it to detect the label of entirely new documents that have never been seen before. In this section, we introduce the computational steps that describe the prediction process, starting with (1) the construction of network training models, (2) the calibration of the algorithm, and (3) the label prediction for each of the ChatGPT-generated articles.